OpenStudy (anonymous):

jacqui has grades of 70 and 79 on her first two tests. if she wants an average of at least 73, what possible scores can she make on her third test?

OpenStudy (anonymous):

\[\frac{70+79+x}{3}\geq 73 \]\[x\geq 70 \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You are welcome.